A box contains 12 similary shaped colored balls let the random variable X represents the number on the ball (numbered from 1-12) and a ball is drawn from a box What is the probobility distrubution of X.What mean and varian of X?







1
Expert's answer
2022-03-04T12:44:55-0500

Since all the 12 similarly shaped colored balls are drawn at random, each of the numbers on the balls has equal chances of being selected, with each having a probability equal to "\frac{1}{12}"


Therefore, the probability distribution of X is:

X P(X)

1 "\frac{1}{12}"

2 "\frac{1}{12}"

3 "\frac{1}{12}"

3 "\frac{1}{12}"

4 "\frac{1}{12}"

5 "\frac{1}{12}"

6 "\frac{1}{12}"

7 "\frac{1}{12}"

8 "\frac{1}{12}"

9 "\frac{1}{12}"

10 "\frac{1}{12}"

11 "\frac{1}{12}"

12 "\frac{1}{12}"


The mean of X

μ = Σ XiPi 

= 1("\frac{1}{12}") + 2("\frac{1}{12}") + 3("\frac{1}{12}") +4("\frac{1}{12}") + 5("\frac{1}{12}") + 6("\frac{1}{12}") + 7("\frac{1}{12}") +8("\frac{1}{12}")+ 9("\frac{1}{12}")

+10("\frac{1}{12}") + 11("\frac{1}{12}") +12("\frac{1}{12}")

= "\frac{78}{12}"

= 6.5

Answer: Mean of X = 6.5

The variance X

σX2 = Σ(x – μ)2⋅ P(x)

="\frac{1}{12}"(1-6.5)2 + "\frac{1}{12}"(2-6.5)2 + "\frac{1}{12}"(3-6.5)2+ "\frac{1}{12}"(4-6.5)2 + "\frac{1}{12}"(5-6.5)2 + "\frac{1}{12}"(6-6.5)2 +

"\frac{1}{12}"(7-6.5)2+ "\frac{1}{12}"(8-6.5)2+ "\frac{1}{12}"(9-6.5)2 + "\frac{1}{12}"(10-6.5)2 + "\frac{1}{12}"(11-6.5)2+ "\frac{1}{12}"(12-6.5)2


=  143/12

= 11.9167

Answer: Variance of X = 11.9167